# SNIC: Synthesized Noisy Images using Calibration

### v1

@@ 4.4 Calibration Procedure
We calibrated noise models for the iPhone 11 Pro main camera, iPhone 11 Pro telephoto camera, Sony RX100 IV, and Sony A7R III. The iPhone main sensor has 12 megapixels and a fixed 35mm equivalent of 26mm. The telephoto camera is also 12 megapixels with a fixed 35mm equivalent of 52 mm. The Sony RX100 IV captures 20 megapixel images with an attached zoom lens with an approximate 35mm equivalent range of 24mm to 70mm. For the 42 megapixel full-frame Sony A7R III, we attached the Sony FE 24-105mm F4 G OSS lens. This was also the lens used for the capture of all calibration and validation images. That lens was chosen to ensure a similar focal length range with the iPhone and RX100 IV when capturing dataset images. For the Sony RX100 IV and A7R III, we captured images at a median aperture and focal length to minimize lens distortion and shading.
@@ 4.4.1 Flat-Field Calibration
Figure 5: Flat-field crop area for Sony A7R III showing low variance in pixel intensity.
After collecting the flat-field images and identifying the crop area, for each ISO, we grouped and sorted the images into pairs by exposure time. The Bayer pattern data was cropped and the channel means were computed. Images that were too dark (mean <= 0.02) or too bright (mean >= 0.98) were discarded to avoid clipped data. We tiled each pair of surviving images. For each tile pair, we computed the tile means and computed the variance by subtracting one tile from its partner. To estimate the Poisson-Gaussian noise parameters, we used a two-stage procedure: first, a HuberRegressor identified and removed outliers from the mean-variance relationship; second, a standard least-squares linear regression on the resulting cleaned data yielded the final Poisson-Gaussian noise model parameters (photon shot noise coefficient $a$ and read noise variance $b$).
@@ 4.4.2 Dark Current Capture
Read noise is extremely difficult to fully characterize in modern sensors [1]. While flat-field calibration can characterize read noise, a more robust technique uses dark frames. Dark frames are captured with a lens cap when possible, and in a dark room or environment. We placed cameras face down inside black fabric pouches and placed those inside sealed boxes, with only a small cut out for the tethering cable. Then per the technique in [1] we captured 10 dark frames at each ISO at a standard exposure time of 1 second. It is essential to disable any in-camera noise reduction to get accurate dark frame images. Those dark frame images are used in the synthesis process as described below.
@@ 4.5 Tuning the Noise Model
After observing non-linearity in the initial noise models for the Sony RX100 IV and the iPhone, we confirmed the issue by visually comparing real noisy images with calibration images, especially at high ISO values where it is most noticeable. To normalize the noise model across ISO settings, we fit a power-law relationship between shot noise parameter $a$ and ISO sensitivity. For each color channel, we constrained the fit to pass exactly through measurements at the 2nd and 4th ISO values (e.g., ISO 50 and 200 for the iPhone), which empirically exhibited high measurement reliability (R² > 0.999). The relationship follows:
@@ 4.6 Noise Synthesis
To generate noisy images, we start with low-ISO RAW images from each camera. These “validation” images have texture and mixed subject matter, rather than uniform, flat-fields. When capturing these low ISO images, we also capture pairs of noisy images at a range of ISO values. All validation images are captured using the same tripod and tethering software used for calibration to minimize motion between captures. For each target noisy ISO, we load the calibrated per-channel noise model and, optionally, the list of dark frame images. Not all cameras benefit from dark current synthesis. Specifically we found the Sony RX100 IV and Sony A7R III noise models were extremely accurate and additional dark current worsened the synthesis quality. For the iPhone cameras, introduction of dark current improved the quality.
@@
When including dark current, we do not use the read noise from the noise model. Instead, one of the ten dark frames is randomly selected. Following [1], rather than random sampling from the chosen dark current frame, we inject dark current with spatial alignment maintained to increase the accuracy of the synthesis. Once the final noisy Bayer data was generated, we produced multiple outputs: First, we wrote the noisy Bayer data to a binary file along with image metadata in a separate JSON file. Second, we demosaiced the noisy Bayer data and rendered them to 16-bit TIFF files. Then, we demosaiced the clean RAW and validation RAW images and rendered them to 16-bit TIFFs using the same RAW decoding parameters. For the dataset, we also created noisy RAW images. For those, we developed a C++ command line tool, linked to the Adobe DNG SDK, to read the binary and JSON files written earlier to create standard RAW DNGs.The result is a collection of RAW and TIFF images: clean, real noise, and synthesized noise.
@@ 5 Results
Our images with noise introduced using calibrated and tuned noise models are more realistic than images using DNG noise models. The improvement in realism is greater for the iPhone cameras and lesser for the Sony RX100 IV and Sony A7R III, but superior in all cases. In addition, we evaluated our physically modeled noise with the NAFNet [12] denoising model, which showed superior comparative results when tested with our images.
@@ 5.1 Evaluating Synthesis Quality
To assess synthesis quality, we compared synthesized noisy images to real noisy images. To establish a baseline, we compared pairs of real noisy images captured in succession at each ISO. We considered three metrics: LPIPS [13] (lower is better), SSIM [14] (higher is better), and PSNR (higher is better). While PSNR would seem to be the obvious metric, it is only suitable for comparing clean images to denoised images. To confirm that, we computed the PSNR for the iPhone 11 Pro main camera, using a set of real noisy image pairs. The PSNR value ranged from -69 dB to below -80 dB. The same issue affects the SSIM metric. When used to compare two real noisy images, the SSIM metric approached 0.0 (no similarity). In contrast, LPIPS, a learned perceptual similarity metric, was a more effective metric. A low LPIPS score indicates greater similarity between two images, and unlike PSNR or SSIM, this metric consistently identified the pairs of real noisy images as similar, rather than dissimilar. When comparing different synthesis techniques, we computed the LPIPS value at each ISO for a real noisy image pair, and then plotted the absolute value of the difference between that value and the LPIPS value of a pair of synthesized noisy image and a real noisy image. We also plotted absolute LPIPS values for real noisy pairs and pairs of synthesized noisy images and real noisy images.
@@ 5.2 LPIPS results
We have provided plots for the four camera sensors, using both absolute and relative measurements. The left plot shows the LPIPS absolute value (lower is better) for each combination of Real Noisy 1, Real Noisy 2, DNG model, and ours. The right plot shows a reduced chart that is relative to the Real Noisy 1 / Real Noisy 2 pair, where closer to zero is best. The calibrated and tuned noise for the two iPhone cameras are clearly superior to the DNG model noise and close to the real noisy images, even at higher ISO values.
(a) Absolute LPIPS
(b) Relative LPIPS
Figure 10: iPhone 11 Pro main camera
Figure 8 shows full-size crops of real noise, DNG model noise, and our tuned and calibrated noise for the iPhone 11 Pro telephoto lens at ISO 1600.
5.3 NAFNet Denoising Results
The Nonlinear Activation Free Network (NAFNet) [12] is a simple denoising model that, despite its simplicity, achieves SOTA results. We used their model pretrained on the real-noise SIDD dataset (NAFNet-SIDD-width64) [6]. We constructed LMDB datasets for each camera with three noise approaches: real noise, the DNG model, and our calibrated and tuned model. Each was paired with the low-ISO ground truth images.
Because the SIDD validation set uses 256×256 pixel patches, we generated 256×256 patches from our full-resolution validation images. Our validation set comprises 79164 patch pairs — over 60× more than SIDD’s 1280 validation patches. The iPhone cameras contributed over 8000 patches each, while higher-resolution Sony cameras produced over 17000 patches each.
This experiment evaluates how well NAFNet generalizes to different noise types (real, DNG model, and ours) using identical scenes, cameras, and ISO values, rather than comparing absolute performance with SIDD. Using PSNR and SSIM metrics, NAFNet denoised our synthesized noisy images significantly more accurately than noisy images created with DNG models. For the iPhone main camera, the PSNR gap between our synthesized noise and real noise was 1.89 dB versus 5.33 dB for DNG model synthesis — a 64.5% improvement. The SSIM gap was 0.083 versus 0.141, a 41.1% improvement. These results demonstrate that our calibrated models produce substantially more realistic noise than DNG models. Complete results appear in Table 2.
@@ 6 Dataset
For each sensor, we captured the same 30 scenes (15 indoor and 15 outdoor). The scenes cover a range of lighting, colors, and textures. All images were captured in RAW format at the sensor’s lowest real ISO. Then, we injected noise for a range of ISO values using our calibrated and tuned noise models into the low-ISO RAW original, using the dark frame database when appropriate. For the dataset, for each camera and scene, we included the low-ISO RAW image and a 16-bit TIFF image rendered from that RAW image. Then, we included both a noisy RAW DNG and a noisy 16-bit TIFF at a range of ISO values. The dataset contains approximately 6600 images. The dataset is located on the Harvard Dataverse at https://doi.org/10.7910/DVN/SGHDCP.
@@ 7 Conclusion
Appendix A Appendix
Following are additional plots of LPIPS results. The “Absolute LPIPS” figures score pairs of images in absolute terms, while the “Relative LPIPS” figures score each synthesis technique against a zero-baseline of the pairs of real noisy images. In all cases for these LPIPS figures, a lower score is better. Following the figures are comparisons of real and synthesized noisy images from each camera.
@@ Appendix A Appendix
Figure 11: iPhone LPIPS measurements
(a) Sony RX100 IV
(b) Sony A7R III
Figure 12: Sony LPIPS measurements

### v2

@@ 3.4 Calibration Procedure
We calibrated noise models for the iPhone 11 Pro main camera, iPhone 11 Pro telephoto camera, Sony RX100 IV, and Sony A7R III. The iPhone main sensor has 12 megapixels and a fixed 35mm equivalent of 26mm. The telephoto camera is also 12 megapixels with a fixed 35mm equivalent of 52 mm. The Sony RX100 IV captures 20 megapixel images with an attached zoom lens with an approximate 35mm equivalent range of 24mm to 70mm. For the 42 megapixel full-frame Sony A7R III, we attached the Sony FE 24-105mm F4 G OSS lens. This was also the lens used for the capture of all calibration, tuning, and dataset images. That lens was chosen to ensure a similar focal length range with the iPhone and RX100 IV when capturing dataset images. For the Sony RX100 IV and A7R III, we captured images at a median aperture and focal length to minimize lens distortion and shading.
@@ 3.4.1 Flat-Field Calibration
Figure 5: Flat-field crop (512x512 pixels) for Sony A7R III showing low variance in pixel intensity.
After collecting the flat-field images and identifying the crop area, for each ISO, we grouped and sorted the images into pairs by exposure time. The Bayer pattern data was cropped and the channel means were computed. Images that were too dark ($\text{mean}\leq 0.02$) or too bright ($\text{mean}\geq 0.98$) were discarded to avoid clipped data. We tiled each pair of surviving images. For each tile pair, we computed the tile means and computed the variance by subtracting one tile from its partner. To estimate the Poisson-Gaussian noise parameters, we used a two-stage procedure: first, a HuberRegressor identified and removed outliers from the mean-variance relationship; second, a standard least-squares linear regression on the resulting cleaned data yielded the final Poisson-Gaussian noise model parameters (photon shot noise coefficient $a$ and read noise variance $b$).
@@ 3.4.2 Dark Current Capture
Read noise is extremely difficult to fully characterize in modern sensors (Zhang et al., 2021). While flat-field calibration can characterize read noise, a more robust technique uses dark frames. Dark frames are captured with a lens cap when possible, and in a dark room or environment. We placed cameras face down inside black fabric pouches and placed those inside sealed boxes, with a small hole cut for the tethering cable. Then per the technique in Zhang et al. (2021), we captured 10 dark frames at each ISO at a standard exposure time of 1 second. We took care to disable any in-camera noise reduction to get accurate dark frame images. Those dark frame images are used in the synthesis process as described below.
@@ 3.5 Tuning the Noise Model
After observing non-linearity in the initial noise models for the Sony RX100 IV and the iPhone, we confirmed the issue by visually comparing real noisy images with calibration images, especially at high ISO values where it is most noticeable. To normalize the noise model across ISO settings, we fit a power-law relationship between shot noise parameter $a$ and ISO sensitivity. For each color channel, we constrained the fit to pass exactly through measurements at the 2nd and 4th ISO values (e.g., ISO 50 and 200 for the iPhone), which empirically exhibited high measurement reliability ($R^{2}>0.999$). The relationship follows:
@@ 3.6 Noise Synthesis
We employed noise synthesis at two different stages. The first was after calibration and calculation of the initial noise models. We captured a series of ”tuning images,” which are sets of clean and real noisy images (described above in Non-linearity in iPhone Noise Models). These tuning images were also used to evaluate synthesis quality, as described in the Results section. Once the noise models were tuned and evaluated, we created the SNIC dataset by adding synthesized noise to a separate set of clean low-ISO dataset images.
For each target noisy ISO, we load the calibrated per-channel (and per-ISO) noise model and, optionally, the list of dark frame images. Not all cameras benefit from dark current synthesis. Specifically, we found the Sony RX100 IV and Sony A7R III noise models were extremely accurate and additional dark current worsened the synthesis quality. For the iPhone cameras, introduction of dark current improved the quality.
@@
When including dark current, we do not use the read noise from the noise model. Instead, one of the ten dark frames is randomly selected. Following Zhang et al. (2021), rather than random sampling from the chosen dark current frame, we inject dark current with spatial alignment maintained to increase the accuracy of the synthesis. Once the final noisy Bayer data was generated, we produced multiple outputs: First, we wrote the noisy Bayer data to a binary file along with image metadata in a separate JSON file. Second, we demosaiced the noisy Bayer data and rendered them to 16-bit TIFF files. Then, we demosaiced both clean and noisy RAW images and rendered each image to a 16-bit TIFF using the same RAW decoding parameters. For the dataset, we also created noisy RAW images from the clean RAWs. For those, we developed a C++ command line tool, linked to the Adobe DNG SDK, to read the binary and JSON files written earlier to create standard RAW DNGs. The result is a collection of clean and synthesized noisy images in both RAW and TIFF formats.
@@ 4 Results
Our images with noise introduced using calibrated and tuned noise models are more realistic than images using DNG noise models. The improvement in realism is greater for the iPhone cameras and lesser for the Sony RX100 IV and Sony A7R III, but superior in all cases. In addition, we evaluated our physically modeled noise with the NAFNet denoising model (Chen et al., 2022), which showed superior comparative results when tested with our images. Figure 8 shows full-size crops of real noise, DNG model noise, and our tuned and calibrated noise for the iPhone 11 Pro telephoto lens at ISO 1600.
@@ 4.1 Evaluating Synthesis Quality
To assess synthesis quality, we compared synthesized noisy images to real noisy images. We used the tuning images from the first noise synthesis stage, and added synthesized noise to the low-ISO tuning images using the final tuned noise models. To establish a baseline, we compared pairs of real noisy images captured in succession at each ISO. We considered three metrics: LPIPS (Zhang et al., 2018) (lower is better), SSIM (Wang et al., 2004) (higher is better), and PSNR (higher is better). While PSNR would seem to be the obvious metric, it is only suitable for comparing clean images to denoised images. To confirm that, we computed the PSNR for the iPhone 11 Pro main camera, using a set of real noisy image pairs. The PSNR value for real noisy image pairs ranged from -69 dB to below -80 dB. The same issue affects the SSIM metric. When used to compare two real noisy images, the SSIM metric approached 0.0 (no similarity). In contrast, LPIPS, a learned perceptual similarity metric, was more effective. A low LPIPS score indicates greater similarity between two images, and unlike PSNR or SSIM, this metric consistently identified the pairs of real noisy images as similar, rather than dissimilar. When comparing different synthesis techniques, we computed the LPIPS value at each ISO for a real noisy image pair, and then plotted the absolute value of the difference between that value and the LPIPS value of a pair of synthesized noisy image and a real noisy image. We also plotted absolute LPIPS values for real noisy pairs and pairs of synthesized noisy images and real noisy images. Figure 9 illustrates this for the iPhone 11 Pro camera.
@@ 4.2 LPIPS results
We have provided plots for the four camera sensors, using both absolute and relative measurements. The left plot shows the LPIPS absolute value (lower is better) for each combination of Real Noisy 1, Real Noisy 2, DNG model, and ours. The right plot shows a reduced chart that is relative to the Real Noisy 1 / Real Noisy 2 pair, where closer to zero is best. The calibrated and tuned noise for the two iPhone cameras are clearly superior to the DNG model noise and close to the real noisy images, even at higher ISO values. Figure 10 shows LPIPS results for the iPhone 11 Pro main camera.
(a) Absolute LPIPS
(b) Relative LPIPS
Figure 10: iPhone 11 Pro main camera
4.3 NAFNet Denoising Results
The Nonlinear Activation Free Network (NAFNet) (Chen et al., 2022) is a denoising model that, despite its simplicity, achieves state-of-the-art results. We used their model pretrained on the real-noise SIDD dataset (NAFNet-SIDD-width64) (Abdelhamed et al., 2018). We constructed LMDB datasets for each camera with three noise approaches: real noise, the DNG model, and our calibrated and tuned model. Each was paired with the low-ISO ground truth images.
[…]
The purpose of this experiment was to evaluate how well NAFNet generalizes to different noise types (real, DNG model, and ours) using identical scenes, cameras, and ISO values, rather than comparing absolute performance with SIDD. Using PSNR and SSIM metrics, NAFNet denoised our synthesized noisy images significantly more accurately than noisy images created with DNG models. For the iPhone main camera, the PSNR gap between our synthesized noise and real noise was 1.89 dB versus 5.33 dB for DNG model synthesis — a 64.5% improvement. The SSIM gap was 0.083 versus 0.141, a 41.1% improvement. These results demonstrate that our calibrated models produce substantially more realistic noise than DNG models.
@@ 5 Dataset
These cameras were selected to span a range of camera types (smartphone, point-and-shoot, and DSLR), a range of sensor sizes and capabilities, and for their tethering capabilities. In addition, the Sony cameras and lenses were chosen because their focal lengths are similar to those of the fixed lens iPhone cameras. Tethering eliminates the need to touch the camera when configuring the camera or capturing images. This is important because touching the camera invariably moves it a small amount, affecting alignment and quality, especially at longer exposure times. The models provided in the SNIC dataset are specific to these cameras. However, the calibration code and techniques are generic and can be used by researchers to create noise models for other cameras.
@@
For each sensor, we captured the same 30 scenes (15 indoor and 15 outdoor). The scenes and images are distinct from the tuning images described in the Noise Synthesis section, and cover a range of locations, lighting, colors, and textures. Further, there are no real noisy images captured for the dataset. All images were captured in RAW format at the sensor’s lowest real ISO. Then, we injected noise for a range of ISO values using our calibrated and tuned noise models into the low-ISO RAW original, using the dark frame database when appropriate. For the dataset, for each camera and scene, we included the low-ISO RAW image and a 16-bit TIFF image rendered from that RAW image. Then, we included both a noisy RAW DNG and a noisy 16-bit TIFF at a range of ISO values. The dataset contains approximately 6600 images. The dataset is located on the Harvard Dataverse at https://doi.org/10.7910/DVN/SGHDCP. A datasheet for the dataset is provided as supplementary material.
@@ 6 Conclusion
Broader Impact Statement
The SNIC dataset can be used both to train and to test noise reduction models. Such models can be used in a wide range of applications, such as image processing and computer vision. The calibration methods described and code provided would allow others to create datasets for other cameras, or to validate existing noise models. A high quality dataset in RAW and TIFF formats allows researchers to work at different spots in either linear or gamma-corrected space, which offers the potential for more sophisticated noise reduction models to be created and validated. The code provides calibration, tuning, and injection scripts to create new high quality datasets. Improved noise reduction models would result in higher quality images and videos, especially in low light environments.
At the same time, noise reduction can be used in computer vision to clean up still images and video frames from video surveillance. While the dataset and code do not themselves provide the ability to enhance surveillance, improved denoising could enhance surveillance capabilities through better quality low-light video frames.
Acknowledgments and Disclosure of Funding
No external financial support or grants were received from any public, commercial, or not-for-profit entities for the research, authorship, or publication of this article. The author declares that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Additional Results
Following are additional plots of LPIPS results. The “Absolute LPIPS” figures score pairs of tuning images in absolute terms, while the “Relative LPIPS” figures score each synthesis technique against a zero-baseline of the pairs of real noisy images. In all cases for these LPIPS figures, a lower score is better. Following the figures are comparisons of real and synthesized noisy images from each camera.
@@ Appendix A Additional Results
(c) Sony RX100 IV
(d) Sony A7R III
Figure 11: LPIPS measurements for all cameras